# State of the Art in Text Classification for South Slavic Languages: Fine-Tuning or Prompting?

### v1

@@ 5.1.  State of the Art in Text Classification Tasks
Figure 1(a) shows that both open-source and closed-source LLMs, used in a zero-shot prompting setup on the sentiment identification task, achieve performance that is comparable or even significantly higher to that of a fine-tuned BERT-like model trained on a large manually-annotated sentiment dataset. The only models that consistently perform worse than the fine-tuned BERT-like model are GPT-3.5-Turbo and DeepSeek-R1-Distill. Sentiment classification appears broad enough that more potent LLMs can interpret label definitions effectively without task-specific fine-tuning, reducing the benefit of additional training.
@@ 5.2.  Comparison of Large Language Models
Figure 2 shows the performance of open-source and closed-source LLMs, used via prompting, on the tasks of sentiment identification, automatic genre identification, news topic classification, and parliamentary topic classification. The DeepSeek-R1-Distill model is not included in the comparison, as it performs significantly worse than other models, as shown in Figure 1.
While different models perform best across different languages and test datasets, a clear trend emerges: the top-performing models across all four benchmark families are the closed-source GPT-4o and GPT-5 from OpenAI, along with Gemini 2.5 Flash. Although GPT-5 is newer and reportedly more powerful, it does not outperform GPT-4o on all benchmarks. Among open-source models, Gemma 3 generally achieves the best results in sentiment identification (Figure 2(a)) and news topic classification (Figure 2(c)). For automatic genre identification (Figure 2(b)) and parliamentary topic classification (Figure 2(d)), rankings of open-source models vary by language. Overall, the weakest performance is observed with the older closed-source GPT-3.5-Turbo model, highlighting the rapid progress in both open-source and closed-source model development.
@@ 5.3.  Advantages and Disadvantages of LLMs
A clear advantage of LLMs is that they do not require manually-annotated training data for specific tasks, yet still achieve strong performance when provided only with task instructions and brief label descriptions. However, these models are significantly more computationally expensive than fine-tuned BERT-like models. While closed-source models deliver the best performance, as shown in previous sections, they come with several limitations: they are costly to use, their architectures and pre-training data are not publicly disclosed, and access through APIs hinders reproducibility, in contrast to open-source LLMs and fine-tuned BERT-like models.
What is more, the inference speed of all LLMs is significantly slower than that of a fine-tuned BERT-like model. As shown in Figure 3, the fine-tuned BERT-like model achieves one of the highest macro-F1 scores on the topic classification task for parliamentary speeches, while maintaining a very low inference time of just 0.02 seconds per instance. In contrast, most LLMs have inference times between 0.6 and 1.4 seconds per instance, making them three to seven times slower for annotating the same dataset. The slowest model, GPT-5, takes 5.5 seconds per instance, which renders it impractical for large-scale automatic annotation of text collections. In this regard, fine-tuned BERT-like models offer a key advantage due to their lower computational cost and higher inference speed. Moreover, they can be trained on training data that is annotated by LLMs using the recently introduced LLM teacher-student paradigm (kuzman-iptc-classification), which considerably reduces the effort needed to develop task-specific models.
@@ 5.4.  Performance on English versus on South Slavic languages
As shown in Figures 2(a) and 2(d), LLMs generally perform worse on Bosnian compared to other languages. However, as shown in Table 3, the differences in macro-F1 scores between English and the average of macro-F1 scores for South Slavic languages are relatively small for sentiment identification, ranging from 2 to 7 points. For topic classification, the performance gap is slightly larger, ranging from 3 to 10 points. This is likely due to the increased difficulty of the task, which involves greater label granularity: 22 labels compared to just 3 in sentiment classification. These findings partially confirm hypothesis H2, which stated that LLMs, when used in a zero-shot setup, perform comparably on text classification tasks in South Slavic languages as they do on English.
Interestingly, even the open-source LLaMA 3.3 model – reported to support only eight languages, excluding the South Slavic group – does not show a substantial performance drop when applied to South Slavic languages compared to English.
@@ 6.  Conclusion
Our results show that LLMs used with prompting, where only a brief task description and labels were provided, achieved strong results across all tasks and languages, particularly the closed-source GPT-4o (openai-gpt4o), GPT-5 (gpt-5) and Gemini 2.5 Flash (comanici2025gemini) models. The performance of LLMs is comparable or higher to that of fine-tuned BERT-like models specialized for the tasks. On the sentiment identification task, most open-source and closed-source LLMs outperformed the fine-tuned model, demonstrating strong general knowledge of the notion of sentiment. For genre and topic classification, however, fine-tuning BERT-like models remain beneficial, as these tasks rely on predefined label sets and fine-tuning aligns the models more closely with the task requirements.
@@
Interestingly, LLMs perform similarly in English and South Slavic languages, with rather minor drops in micro- and macro-F1 scores, namely a drop of 2 to 7 points in terms of macro-F1 scores on sentiment classification, and a slightly higher drop from 3 to 10 points on topic classification in parliamentary texts. This suggests that the gap in multilingual performance is smaller than expected, even for open-source models not explicitly dedicated to these languages.
@@
Although large language models offer impressive zero-shot performance and reduce the need for annotated data, they come with higher computational costs and are more prone to producing invalid labels. Moreover, their inference speed is at least three times slower than that of the fine-tuned BERT-like models. Thus, their use in use cases with extensive data to be processed, such as automatic enrichment of large corpora with text categories, remains impractical due to their high computational demands. In contrast, fine-tuned BERT-like models are more computationally efficient and can be better tailored to the specific characteristics of a task and its domain. They remain a practical and reliable choice for text classification tasks, especially when computational resources are limited, high inference speed is desired or output reliability is critical. Moreover, it is possible to combine the strengths of both approaches, as proposed by the LLM teacher-student pipeline paradigm (kuzman-iptc-classification): LLMs can be used to automatically annotate training data, reducing the need for costly and time-consuming manual annotation, while fine-tuned BERT-like models can then be trained on these datasets.
@@
This study represents only an initial step to systematically benchmark text classification performance in South Slavic languages. Although our evaluation includes four diverse benchmark families, some of the test datasets remain relatively small. Future work will aim to increase dataset sizes, include more South Slavic languages and dialects, and introduce additional classification tasks. As new large language models continue to emerge rapidly, it will also be important to establish ongoing evaluations to track whether their performance continues to improve, particularly on South Slavic languages. Importantly, this study only evaluated the performance of LLMs in a zero-shot prompting setup. In future work, we plan to extend the evaluation to include few-shot prompting and fine-tuning on training data. To support further research and facilitate reproducibility, we have made all code, evaluation scripts, and results publicly available. [^5] https://github.com/TajaKuzman/Benchmarking-Text-Classification-on-South-Slavic
@@ 8.  Acknowledgements
We would like to thank the developers of the llm.ijs.si service (maric_local_llms) for establishing the LLM inference platform deployed at the Jožef Stefan Institute, which provided convenient access to the open-source large language models used in this study. We also thank the annotators of the test datasets for their diligence and the time devoted to manual annotation, which resulted in the high-quality evaluation datasets used in this work. Lastly, we would like to thank the CLASSLA knowledge centre for South Slavic languages and the Slovenian CLARIN.SI infrastructure for their valuable support.
@@ A.1.  Benchmarking Datasets
include Croatian, Serbian, Bosnian, and English data from the multilingual sentiment dataset of parliamentary debates ParlaSent 1.0 (mochtak2024parlasent, parlasent-repository). [^6] Available in the CLARIN.SI repository at http://hdl.handle.net/11356/1585 and in the Hugging Face repository at https://huggingface.co/datasets/classla/ParlaSent. The dataset comprises sentences that were randomly sampled from Croatian, Serbian, Bosnian and British parliamentary corpora and manually annotated with reported inter-annotator agreement ranging from 0.53 to 0.66 in Krippendorff’s alpha (krippendorff2018content). The annotation involved a more granular six-level sentiment polarity scale that has been mapped to a three-level sentiment polarity scale which we use in our experiments: negative (0), neutral (1), and positive (2).
@@
comprise the English EN-GINCO dataset (kuzman2023automatic) and a multilingual X-GINCO dataset from the AGILE benchmark for Automatic Genre Identification. [^7] https://github.com/TajaKuzman/AGILE-Automatic-Genre-Identification-Benchmark The test instances were sampled from the enTenTen20 English web corpus (jakubivcek2013tenten) and the MaCoCu multilingual web corpus collection (banon2022macocu). They were manually annotated by experts with a background in linguistics and computational linguistics who had experience with previous genre annotation campaigns (kuzman-rupnik-ljubei:2022:LREC; kuzman2023automatic) where they reached an acceptable inter-annotator agreement of 0.71 in nominal Krippendorff’s alpha (krippendorff2018content). While the X-GINCO dataset comprises numerous European languages, for the purposes of this study, we focus on three South Slavic languages: Croatian, Macedonian, and Slovenian. The test datasets use the X-GENRE annotation schema (kuzman2023automatic) that includes the following genre labels: Information/Explanation, News, Instruction, Opinion/Argumentation, Forum, Prose/Lyrical, Legal and Promotion. While EN-GINCO and X-GINCO datasets have been annotated by the same annotator with the same schema, one should note that there are important differences between them in terms of their construction – the English test dataset was sampled randomly from the web corpus, resulting in an unbalanced label distribution, while the X-GINCO datasets were curated with more deliberate interventions to ensure a balanced label distribution and a more controlled sampling process. Consequently, the X-GINCO datasets comprise fewer ambiguous instances and could be regarded as an easier test dataset.
@@
(kuzman-iptc-classification) comprise Croatian and Slovenian news articles extracted from the MaCoCu-Genre web corpus collection (macocu-genre) and manually annotated by one annotator. The reliability of the annotator was confirmed on a sample of data that was annotated by an additional annotator. The two annotators reached an acceptable inter-annotator agreement of 0.73 in nominal Krippendorff’s alpha (krippendorff2018content). Text instances are annotated with 17 topic labels from the top level of the IPTC NewsCodes Media Topic hierarchical schema, developed by the International Press Telecommunications Council (IPTC) (iptcGroupsNewsCodes). The datasets are more or less balanced by labels.
@@
comprise parliamentary speeches in Bosnian, Croatian, English, and Serbian, sourced from the ParlaMint 4.1 dataset (parlamint_41; erjavec2025parlamint). These speeches were annotated by a single expert annotator using the 21 CAP categories from the official CAP schema (baumgartner2019comparative), along with an additional Other label. The datasets are approximately balanced across labels. To assess the annotation quality, the Croatian dataset was independently annotated by two additional annotators. Inter-annotator agreement between the expert annotator and the others ranged from 0.62 to 0.68 in Krippendorff’s alpha, which is around the threshold of 0.67 typically considered acceptable for annotation reliability (krippendorff2018content).
@@ A.2.  Models
In the following subsections, we outline the models included in the evaluation – the fine-tuned BERT-like classifiers (Section A.2.1) and the open-source and closed-source LLMs (Section A.2.2).
@@ A.2.1.  Fine-Tuned BERT-like Models
BERT (bidirectional encoder representations from transformers) deep neural models (bert) have revolutionized the field of natural language processing (NLP), outperforming the non-neural methods across various NLP tasks. They have a more complex and computationally expensive architecture featuring transformers – neural networks with self-attention mechanisms (vaswani2017attention) – that significantly improves the efficiency of training models on massive text data. Similarly to decoder-only transformer models, BERT models are pretrained on massive amounts of texts, possibly in multiple languages, which establishes their ability to encode the words and texts in high-dimensional vector spaces (minaee2020deep) and enables their application even across languages in a zero-shot classification scenario. To develop BERT-based classifiers, the pretrained models are trained, that is, fine-tuned, on a training dataset comprising text instances annotated with labels. In our study, we evaluate openly-accessible multilingual fine-tuned BERT-like models that have been already developed in recent related research. Namely, we evaluate the following models:
@@
  • IPTC News Topic classifier [^8] The IPTC News Topic classifier is available in the Hugging Face repository at https://huggingface.co/classla/multilingual-IPTC-news-topic-classifier. (kuzman-iptc-classification) is a multilingual fine-tuned BERT-like model for news topic classification according to the top-level IPTC NewsCodes schema (iptcGroupsNewsCodes). The model is based on the large-sized XLM-RoBERTa model (conneau2020unsupervised) and was fine-tuned on 15,000 training text instances from the EMMediaTopic [^9] The EMMediaTopic training dataset is available in the CLARIN.SI repository at http://hdl.handle.net/11356/1991. dataset (emmediatopic). The training dataset contains news article instances in four languages: Catalan, Croatian, Greek, and Slovenian. The training dataset was annotated using an LLM that was shown to achieve annotation reliability comparable to that of human annotators (kuzman-iptc-classification). This approach is based on the novel methodology that uses the LLM teacher-student pipeline to develop BERT-like classifiers in the absence of manually-annotated training data.
@@
  • XLM-R-ParlaSent (parlasent-model; mochtak2024parlasent) is a domain-specific multilingual transformer model for sentiment identification in parliamentary texts. It is based on the XLM-R-parla pretrained model (xlm-r-parla) that was developed by additionally pretraining the large-sized XLM-RoBERTa model (conneau2020unsupervised) on 1.72 billion words from parliamentary proceedings in 30 European languages. To develop the XLM-R-ParlaSent model, [^10] The XLM-R-ParlaSent model is accessible in the Hugging Face repository at https://huggingface.co/classla/xlm-r-parlasent. the pretrained XLM-R-Parla model was fine-tuned on the ParlaSent sentiment training dataset (mochtak2024parlasent; parlasent-repository) in seven European languages (Bosnian, Croatian, Czech, English, Serbian, Slovak, and Slovenian). The training dataset [^11] The ParlaSent training and test datasets are freely available in the CLARIN.SI repository at http://hdl.handle.net/11356/1868. comprises 13,000 instances sampled from parliamentary proceedings and manually annotated with sentiment labels.
@@
  • ParlaCAP classifier [^12] The ParlaCAP topic classifier is available in the Hugging Face repository at https://huggingface.co/classla/ParlaCAP-Topic-Classifier. (parlacap_model) is a domain-specific multilingual transformer model for topic classification in parliamentary texts based on the CAP schema (baumgartner2019comparative). As the XLM-R-ParlaSent model, this model is based on the XLM-R-parla pretrained model (xlm-r-parla; mochtak2024parlasent). The XLM-R-parla model was then fine-tuned on around 30 thousand speeches from parliamentary debates from the ParlaMint 4.1 parliamentary datasets (parlamint_41; erjavec2025parlamint) in 29 European languages. The training dataset was annotated with the CAP categories by a GPT-4o (openai-gpt4o) model used in a zero-shot prompting fashion, following the LLM teacher-student framework (kuzman-iptc-classification). Based on the inter-annotator agreement, calculated on a sample that was annotated by three human annotators and the LLM annotator, the agreement between the LLM and the human annotators was comparable to the agreement between the human annotators themselves. This indicates that the LLM performs as reliably as human annotators on this task, supporting its use for annotating the training data.
@@
  • X-GENRE classifier (kuzman2023automatic; x-genre-repository) is a multilingual fine-tuned BERT-like model for automatic genre identification. [^13] The X-GENRE classifier is freely available in the Hugging Face repository at https://doi.org/10.57967/hf/0927 and the CLARIN.SI repository at http://hdl.handle.net/11356/1961. The model is based on the base-sized XLM-RoBERTa model (conneau2020unsupervised) and was fine-tuned on the training split of the X-GENRE dataset (x-genre-dataset), which contains 1,772 text instances in Slovenian and English, manually-annotated with genre labels from the X-GENRE schema (kuzman2023automatic).

### v2

@@ 5.1.  State of the Art in Text Classification Tasks
Figure 1(a) shows that both open-weight and closed-source LLMs, used in a zero-shot prompting setup on the sentiment identification task, achieve performance that is comparable or even significantly higher to that of a fine-tuned BERT-like model trained on a large manually-annotated sentiment dataset. The only models that consistently perform worse than the fine-tuned BERT-like model are GPT-3.5-Turbo and DeepSeek-R1-Distill. Sentiment classification appears broad enough that more potent LLMs can interpret label definitions effectively without task-specific fine-tuning, reducing the benefit of additional training.
@@ 5.2.  Comparison of Large Language Models
Figure 2 shows the performance of open-weight and closed-source LLMs, used via prompting, on the tasks of sentiment identification, automatic genre identification, news topic classification, and parliamentary topic classification. The DeepSeek-R1-Distill model is not included in the comparison, as it performs significantly worse than other models, as shown in Figure 1.
While different models perform best across different languages and test datasets, a clear trend emerges: the top-performing models across all four benchmark families are the closed-source GPT-4o and GPT-5 from OpenAI, along with Gemini 2.5 Flash. Although GPT-5 is newer and reportedly more powerful, it does not outperform GPT-4o on all benchmarks. Among open-weight models, Gemma 3 generally achieves the best results in sentiment identification (Figure 2(a)) and news topic classification (Figure 2(c)). For automatic genre identification (Figure 2(b)) and parliamentary topic classification (Figure 2(d)), rankings of open-weight models vary by language. Overall, the weakest performance is observed with the older closed-source GPT-3.5-Turbo model, highlighting the rapid progress in both open-weight and closed-source model development.
@@ 5.3.  Advantages and Disadvantages of LLMs
A clear advantage of LLMs is that they do not require manually-annotated training data for specific tasks, yet still achieve strong performance when provided only with task instructions and brief label descriptions. However, these models are significantly more computationally expensive than fine-tuned BERT-like models. While closed-source models deliver the best performance, as shown in previous sections, they come with several limitations: they are costly to use, their architectures and pre-training data are not publicly disclosed, and access through APIs hinders reproducibility, in contrast to open-weight LLMs and fine-tuned BERT-like models.
What is more, the inference speed of all LLMs is significantly slower than that of a fine-tuned BERT-like model. As shown in Figure 3, the fine-tuned BERT-like model achieves one of the highest macro-F1 scores on the topic classification task for parliamentary speeches, while maintaining a very low inference time of just 0.02 seconds per instance. In contrast, most LLMs have inference times between 0.6 and 1.4 seconds per instance, making them three to seven times slower for annotating the same dataset. The slowest model, GPT-5, takes 5.5 seconds per instance, which renders it impractical for large-scale automatic annotation of text collections. In this regard, fine-tuned BERT-like models offer a key advantage due to their lower computational cost and higher inference speed. Moreover, they can be trained on training data that is annotated by LLMs using the recently introduced LLM teacher-student paradigm (Kuzman and Ljubešić, 2025), which considerably reduces the effort needed to develop task-specific models.
@@ 5.4.  Performance on English versus on South Slavic languages
As shown in Table 3, the differences in macro-F1 scores between English and the average of macro-F1 scores for South Slavic languages are relatively small for sentiment identification, ranging from 2 to 7 points. For topic classification, the performance gap is slightly larger, ranging from 3 to 10 points. This is likely due to the increased difficulty of the task, which involves greater label granularity: 22 labels compared to just 3 in sentiment classification. These findings partially confirm hypothesis H2, which stated that LLMs, when used in a zero-shot setup, perform comparably on text classification tasks in South Slavic languages as they do on English.
Interestingly, even the open-weight LLaMA 3.3 model – reported to support only eight languages, excluding the South Slavic group – does not show a substantial performance drop when applied to South Slavic languages compared to English.
@@ 6.  Conclusion
Our results show that LLMs used with prompting, where only a brief task description and labels were provided, achieved strong results across all tasks and languages, particularly the closed-source GPT-4o (OpenAI, 2024), GPT-5 (OpenAI, 2025) and Gemini 2.5 Flash (Comanici et al., 2025) models. The performance of LLMs is comparable or higher to that of fine-tuned BERT-like models specialized for the tasks. On the sentiment identification task, most open-weight and closed-source LLMs outperformed the fine-tuned model, demonstrating strong general knowledge of the notion of sentiment. For genre and topic classification, however, fine-tuning BERT-like models remains beneficial, as these tasks rely on predefined label sets and fine-tuning aligns the models more closely with the task requirements.
@@
Interestingly, LLMs perform similarly in English and South Slavic languages, with rather minor drops in micro- and macro-F1 scores, namely a drop of 2 to 7 points in terms of macro-F1 scores on sentiment classification, and a slightly higher drop from 3 to 10 points on topic classification in parliamentary texts. This suggests that the gap in multilingual performance is smaller than expected, even for open-weight models not explicitly dedicated to these languages.
@@
Although large language models offer impressive zero-shot performance and reduce the need for annotated data, they come with higher computational costs and are more prone to producing invalid labels. Moreover, their inference speed is at least three times slower than that of the fine-tuned BERT-like models. Thus, their use in use cases with extensive data to be processed, such as automatic enrichment of large corpora with text categories, remains impractical due to their high computational demands. In contrast, fine-tuned BERT-like models are more computationally efficient and can be better tailored to the specific characteristics of a task and its domain. They remain a practical and reliable choice for text classification tasks, especially when computational resources are limited, high inference speed is desired or output reliability is critical. Moreover, it is possible to combine the strengths of both approaches, as proposed by the LLM teacher-student paradigm (Kuzman and Ljubešić, 2025): LLMs can be used to automatically annotate training data, reducing the need for costly and time-consuming manual annotation, while fine-tuned BERT-like models can then be trained on these datasets.
@@
This study represents only an initial step to systematically benchmark text classification performance in South Slavic languages. Although our evaluation includes four diverse benchmark families, some of the test datasets remain relatively small. Future work will aim to increase dataset sizes, include more South Slavic languages and dialects, and introduce additional classification tasks. As new large language models continue to emerge rapidly, it will also be important to establish ongoing evaluations to track whether their performance continues to improve, particularly on South Slavic languages. Importantly, this study only evaluated the performance of LLMs in a zero-shot prompting setup. In future work, we plan to extend the evaluation to include few-shot prompting and fine-tuning on training data. To support further research and facilitate reproducibility, we have made all code, evaluation scripts, and results publicly available. [^5] https://github.com/TajaKuzman/Benchmarking-Text-Classification-on-South-Slavic Additionally, we have developed an interactive dashboard [^6] See the CLASSLA LLM Evaluation Dashboard for South Slavic Languages at https://www.clarin.si/classla-llm-dashboard/. that enables users to explore the results of our evaluation of large language models on the tasks presented in this paper, as well as on additional commonsense reasoning tasks. The dashboard is an ongoing project that monitors the performance of newly released large language models on South Slavic languages and dialects. In future work, we plan to expand both the range of tasks and the coverage of South Slavic languages and dialects included in the dashboard.
@@ 8.  Acknowledgements
We would like to thank the developers of the llm.ijs.si service (Marić et al., 2025) for establishing the LLM inference platform deployed at the Jožef Stefan Institute, which provided convenient access to the open-weight large language models used in this study. We also thank the annotators of the test datasets for their diligence and the time devoted to manual annotation, which resulted in the high-quality evaluation datasets used in this work. Lastly, we would like to thank the CLASSLA knowledge centre for South Slavic languages and the Slovenian CLARIN.SI infrastructure for their valuable support.
@@ A.1.  Benchmarking Datasets
include Croatian, Serbian, Bosnian, and English data from the multilingual sentiment dataset of parliamentary debates ParlaSent 1.0 (Mochtak et al., 2024, Mochtak et al., 2023). [^7] Available in the CLARIN.SI repository at http://hdl.handle.net/11356/1585 and in the Hugging Face repository at https://huggingface.co/datasets/classla/ParlaSent. The dataset comprises sentences that were randomly sampled from Croatian, Serbian, Bosnian and British parliamentary corpora and manually annotated with reported inter-annotator agreement ranging from 0.53 to 0.66 in Krippendorff’s alpha (Krippendorff, 2018). The annotation involved a more granular six-level sentiment polarity scale that has been mapped to a three-level sentiment polarity scale which we use in our experiments: negative (0), neutral (1), and positive (2).
@@
comprise the English EN-GINCO dataset (Kuzman et al., 2023) and a multilingual X-GINCO dataset from the AGILE benchmark for Automatic Genre Identification. [^8] https://github.com/TajaKuzman/AGILE-Automatic-Genre-Identification-Benchmark The test instances were sampled from the enTenTen20 English web corpus (Jakubíček et al., 2013) and the MaCoCu multilingual web corpus collection (Bañón et al., 2022). They were manually annotated by experts with a background in linguistics and computational linguistics who had experience with previous genre annotation campaigns (Kuzman et al., 2022, 2023) where they reached an acceptable inter-annotator agreement of 0.71 in nominal Krippendorff’s alpha (Krippendorff, 2018). While the X-GINCO dataset comprises numerous European languages, for the purposes of this study, we focus on three South Slavic languages: Croatian, Macedonian, and Slovenian. The test datasets use the X-GENRE annotation schema (Kuzman et al., 2023) that includes the following genre labels: Information/Explanation, News, Instruction, Opinion/Argumentation, Forum, Prose/Lyrical, Legal and Promotion. While EN-GINCO and X-GINCO datasets have been annotated by the same annotator with the same schema, one should note that there are important differences between them in terms of their construction – the English test dataset was sampled randomly from the web corpus, resulting in an unbalanced label distribution, while the X-GINCO datasets were curated with more deliberate interventions to ensure a balanced label distribution and a more controlled sampling process. Consequently, the X-GINCO datasets comprise fewer ambiguous instances and could be regarded as an easier test dataset.
@@
(Kuzman and Ljubešić, 2025) comprise Croatian and Slovenian news articles extracted from the MaCoCu-Genre web corpus collection (Kuzman and Ljubešić, 2024b) and manually annotated by one annotator. The reliability of the annotator was confirmed on a sample of data that was annotated by an additional annotator. The two annotators reached an acceptable inter-annotator agreement of 0.73 in nominal Krippendorff’s alpha (Krippendorff, 2018). Text instances are annotated with 17 topic labels from the top level of the IPTC NewsCodes Media Topic hierarchical schema, developed by the International Press Telecommunications Council (IPTC) (IPTC, 2022). The datasets are more or less balanced by labels.
@@
(Kuzman Pungeršek et al., 2026) comprise parliamentary speeches in Bosnian, Croatian, English, and Serbian, sourced from the ParlaMint 4.1 dataset (Erjavec et al., 2024; Erjavec et al., 2025). These speeches were annotated by a single expert annotator using the 21 CAP categories from the official CAP schema (Baumgartner et al., 2019), along with an additional Other label. The datasets are approximately balanced across labels. To assess the annotation quality, the Croatian dataset was independently annotated by two additional annotators. Inter-annotator agreement between the expert annotator and the others ranged from 0.62 to 0.68 in Krippendorff’s alpha, which is around the threshold of 0.67 typically considered acceptable for annotation reliability (Krippendorff, 2018).
@@ A.2.  Models
In the following subsections, we outline the models included in the evaluation – the fine-tuned BERT-like classifiers (Section A.2.1) and the open-weight and closed-source LLMs (Section A.2.2).
@@ A.2.1.  Fine-Tuned BERT-like Models
BERT (bidirectional encoder representations from transformers) deep neural models (Kenton and Toutanova, 2019) have revolutionized the field of natural language processing (NLP), outperforming the non-neural methods across various NLP tasks. They have a more complex and computationally expensive architecture featuring transformers – neural networks with self-attention mechanisms (Vaswani et al., 2017) – that significantly improves the efficiency of training models on massive text data. Similarly to decoder-only transformer models, BERT models are pretrained on massive amounts of texts, possibly in multiple languages, which establishes their ability to encode the words and texts in high-dimensional vector spaces (Minaee et al., 2020) and enables their application even across languages in a zero-shot classification scenario. To develop BERT-based classifiers, the pretrained models are trained, that is, fine-tuned, on a training dataset comprising text instances annotated with labels. In our study, we evaluate openly-accessible multilingual fine-tuned BERT-like models that have been already developed in recent related research. Namely, we evaluate the following models:
@@
  • IPTC News Topic classifier [^9] The IPTC News Topic classifier is available in the Hugging Face repository at https://huggingface.co/classla/multilingual-IPTC-news-topic-classifier. (Kuzman and Ljubešić, 2025) is a multilingual fine-tuned BERT-like model for news topic classification according to the top-level IPTC NewsCodes schema (IPTC, 2022). The model is based on the large-sized XLM-RoBERTa model (Conneau et al., 2020) and was fine-tuned on 15,000 training text instances from the EMMediaTopic [^10] The EMMediaTopic training dataset is available in the CLARIN.SI repository at http://hdl.handle.net/11356/1991. dataset (Kuzman and Ljubešić, 2024c). The training dataset contains news article instances in four languages: Catalan, Croatian, Greek, and Slovenian. The training dataset was annotated using an LLM that was shown to achieve annotation reliability comparable to that of human annotators (Kuzman and Ljubešić, 2025). This approach is based on the novel methodology that uses the LLM teacher-student pipeline to develop BERT-like classifiers in the absence of manually-annotated training data.
@@
  • XLM-R-ParlaSent (Rupnik et al., 2023; Mochtak et al., 2024) is a domain-specific multilingual transformer model for sentiment identification in parliamentary texts. It is based on the XLM-R-parla pretrained model (Ljubešić et al., 2023) that was developed by additionally pretraining the large-sized XLM-RoBERTa model (Conneau et al., 2020) on 1.72 billion words from parliamentary proceedings in 30 European languages. To develop the XLM-R-ParlaSent model, [^11] The XLM-R-ParlaSent model is accessible in the Hugging Face repository at https://huggingface.co/classla/xlm-r-parlasent. the pretrained XLM-R-Parla model was fine-tuned on the ParlaSent sentiment training dataset (Mochtak et al., 2024; Mochtak et al., 2023) in seven European languages (Bosnian, Croatian, Czech, English, Serbian, Slovak, and Slovenian). The training dataset [^12] The ParlaSent training and test datasets are freely available in the CLARIN.SI repository at http://hdl.handle.net/11356/1868. comprises 13,000 instances sampled from parliamentary proceedings and manually annotated with sentiment labels.
@@
  • ParlaCAP classifier [^13] The ParlaCAP topic classifier is available in the Hugging Face repository at https://huggingface.co/classla/ParlaCAP-Topic-Classifier. (Kuzman Pungeršek and Ljubešić, 2025; Kuzman Pungeršek et al., 2026) is a domain-specific multilingual transformer model for topic classification in parliamentary texts based on the CAP schema (Baumgartner et al., 2019). As the XLM-R-ParlaSent model, this model is based on the XLM-R-parla pretrained model (Ljubešić et al., 2023; Mochtak et al., 2024). The XLM-R-parla model was then fine-tuned on the ParlaCAP-train dataset [^14] The ParlaCAP-train training dataset is available in the CLARIN.SI repository at http://hdl.handle.net/11356/2093. (Kuzman Pungeršek and Ljubešić, 2026; Kuzman Pungeršek et al., 2026). The training dataset comprises around 30 thousand speeches from parliamentary debates from the ParlaMint 4.1 parliamentary datasets (Erjavec et al., 2024; Erjavec et al., 2025) in 29 European languages. The training dataset was annotated with the CAP categories by a GPT-4o (OpenAI, 2024) model used in a zero-shot prompting fashion, following the LLM teacher-student framework (Kuzman and Ljubešić, 2025). Based on the inter-annotator agreement, calculated on a sample that was annotated by three human annotators and the LLM annotator, the agreement between the LLM and the human annotators was comparable to the agreement between the human annotators themselves. This indicates that the LLM performs as reliably as human annotators on this task, supporting its use for annotating the training data.
@@
  • X-GENRE classifier (Kuzman et al., 2023; Kuzman and Ljubešić, 2024d) is a multilingual fine-tuned BERT-like model for automatic genre identification. [^15] The X-GENRE classifier is freely available in the Hugging Face repository at https://doi.org/10.57967/hf/0927 and the CLARIN.SI repository at http://hdl.handle.net/11356/1961. The model is based on the base-sized XLM-RoBERTa model (Conneau et al., 2020) and was fine-tuned on the training split of the X-GENRE dataset (Kuzman and Ljubešić, 2024a), which contains 1,772 text instances in Slovenian and English, manually-annotated with genre labels from the X-GENRE schema (Kuzman et al., 2023).